OpenStudy (anonymous):

Please help! I'm solving a system of equations but I can't figure out what I'm doing wrong... The system is: 3x - 6y = 12 -2x + 3y = 6 I distributed 2 to the second equation leaving me with -4x + 6y = 12 so then I got 3x - 6y = 12 -4x + 6y = 12 3x = 12 -4x = 12 -x = 0 x = 0 -------- 3 (0) - 6y = 12 -6y = 12 y = -2 but my potential answers are 14, -14, 24 and -24

OpenStudy (chenping):

try to use substition instead :)

OpenStudy (chenping):

eq1 can be simplified into x=4+2y

OpenStudy (anonymous):

what did you do to get that?

OpenStudy (chenping):

eq 1 divide by 3? :)

OpenStudy (anonymous):

I'll try that. I've never done substitution by dividing an equation before

OpenStudy (phi):

this is good: 3x - 6y = 12 -4x + 6y = 12 add the equations to get -1x + 0 = 24 or -x = 24 x= -24

OpenStudy (phi):

I think you confused adding /subtracting the equations but you want to add because -6y + 6y = 0 (which is what you want)

OpenStudy (chenping):

haha, yea, didnt notice that small mistake xD

OpenStudy (anonymous):

ahhh okay, thak you so much! I appriciate it!

OpenStudy (anonymous):

so does y = -14?

OpenStudy (phi):

yes

OpenStudy (phi):

you can check you answers by putting x=-24 and y= -14 into each equation and seeing if they work. I just did, and both work

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

yw